# Sparse Portfolio Selection via Topological Data Analysis based Clustering

### v1

@@ 5.3 Strategy 2: Sparse Mean-variance and Global Minimum Variance Portfolios
Panel A in Table 9 shows the out-of-sample performance of the TDA-based sparse mean-variance portfolios. We observe that the portfolio $MV2$ achieves the maximum values for mean, SR, and CEQ, and the minimum value of TR, thus outperforming the other portfolios, including TDA-based portfolios and the benchmark portfolios, namely, $MVA$, the naïve strategy and the benchmark index. For instance, $MV2$ achieves the daily Sharpe ratio of .0733 (.3359 per month / 1.1636 per annum) compared to 0.688 and 0.521 achieved by $MVA$ and $SP500$. Further, the least value of the turnover ratio for $MV2$ indicates less re-balancing across the windows and the formation of stable portfolios. In contrast to the performance of TDA-based index-tracking portfolios, $MV4$ fails to consistently beat other TDA-based portfolios as well as the benchmark portfolios. Nonetheless, these findings confirm our conclusion that the TDA-based selection strategy, in particular, based on kernel $K_{2}$, outperforms the benchmark models, thus providing an excellent asset selection criteria. [^19] We also observe that the $MVA$ portfolio has a higher out-of-sample mean, SR, and CEQ compared to those from the naïve $1/N$ portfolio. We also perform hypothesis tests to assess whether the values from $MVA$ and naïve strategy are statistically significant. We observed that the mean and CEQ from $MVA$ is significantly greater than that from naïve strategy at 99% confidence level while the Sharpe ratios from the two are not significantly different, a similar conclusion was drawn in DeMiguel et al. (2009b). Panel B in Table 9 shows the out-of-sample performance of the TDA-based sparse global minimum variance portfolios. A similar conclusion can be drawn that in a $K_{2}$ based portfolio, $GV2$ outperforms other TDA-based portfolios, with the exception that $GV5$ has a lower turnover ratio than $GV2$, but the difference is not very large. Further, $GV2$ beats the benchmark portfolios $GVA$, the naïve, and the index in terms of mean and Sharpe ratio. These findings were further confirmed using the statistical tests for mean, SR, and CEQ; for instance, we observed that $MV2$ has a significantly higher mean, SR, and CEQ than $MVA$ portfolios at a 95% confidence level.
@@ B. Global Minimum Variance Portfolios
COVID-19 Robustness Check Next, we analyze the portfolio performance in the more volatile period from July 2019 to mid-November 2020, which includes the COVID-19 pandemic. Table 11 presents the out-of-sample performance. Panel A clearly shows that not all sparse TDA-based portfolios perform well. This observation is not entirely surprising because of the increased market fluctuations during the pandemic. Nevertheless, the TDA-based portfolio $MV_{2}$ outperforms the $MVA$, the naive portfolio, and the benchmark index. This is demonstrated with significant gaps; for instance, the Sharpe ratio of 4.63E-02 achieved by $MV_{2}$ is significantly better (almost double) than 2.24E-02 and 2.86E-02 achieved by $MVA$ and the naive portfolio, respectively. Unlike TDA-based mean-variance portfolios, there is no clear pattern in the performance of TDA-based global minimum variance portfolios. Nonetheless, these portfolios, in particular, $GV_{2}$ outperform $GVA$ and naïve portfolio, thus confirming our findings that TDA-based portfolios, in particular, based on the kernel $K_{2}$ are not only sparse but also perform superior to traditional models even in turbulent market conditions, safeguarding against significant losses. These results underline the out-of-sample efficacy of TDA-based models in managing risk and delivering returns.
@@
Mean 7.40E-04 5.68E-04 6.27E-04 8.25E-04 3.86E-04 2.90E-05 1.78E-04 6.14E-04 6.73E-04
SR 6.16E-02 3.01E-02 5.25E-02 4.73E-02 2.12E-02 1.65E-03 1.53E-02 2.86E-02 3.49E-02
CEQ 6.68E-04 3.90E-04 5.55E-04 6.73E-04 2.20E-04 -1.25E-04 1.10E-04 3.84E-04 4.87E-04
TR 1.54E+00 1.91E+00 1.47E+00 1.94E+00 1.74E+00 1.64E+00
@@ 6 Conclusion
Appendix
6.1 Case study: Time series clustering
@@
In this section, we demonstrate the superior performance of the newly introduced distance measures by applying clustering on a standard data set from the machine learning literature. More precisely, we show that the distance measures $AWD$ and $DWD$ that quantify the dissimilarity between persistence diagrams ${\mathcal{D}}_{x}$ and ${\mathcal{D}}_{y}$ outperform the conventional Wasserstein distance ($WD$). Furthermore, we show that the proposed distance measure on landscapes $ALD$ and $DLD$ outperform the traditional distance between landscapes ($LD$), defined as $||\lambda_{x}-\lambda_{y}||_{p}$ where $\lambda_{x}$ and $\lambda_{y}$ are landscapes corresponding to the PD ${\mathcal{D}}_{x}$ and ${\mathcal{D}}_{y}$. For comparison purposes, we also compare the traditional distances, denoted by $d_{1}$ and $d_{2}$, defined in terms of Spearman’s rank correlation $\rho_{1}$ and Pearson’s correlation coefficient $\rho_{2}$, as follows $d_{i}(x,y)=\sqrt{2(1-\rho_{i}(x,y))}~{}~{}i\in\{1,2\}.$ Finally, for completeness, we also consider the Euclidean distance $ED$. We evaluate the clustering performance in terms of the accuracy ($ACC$) He et al. (2019), which ranges from 0 to 1, with a higher value indicating better clustering results.
@@
We collect the data set from the UCI machine learning repository that contains 600 examples of control charts with 60 instances each, synthetically generated by the process in Alcock et al. (1999). The data consists of six different classes of control charts: Normal and cyclic, increasing and decreasing trend, upward and downward shift.
@@
Table 12: Accuracy of clustering results with different distance metrics. Each value $v$ is to be read as $v\times 10^{-2}$
@@
Table 12 presents the clustering results based on two standard algorithms, the affinity propagation clustering (APC) with $K$ clusters and the $K$-medoids with the distance measures mentioned above. [^20] We use Gaussian exponential kernel as similarity measure defined using these distance measures as an input for APC. We fix embedding dimension $d=2$, and we take $\tau=1$ with $p=1$ for calculating the TDA-based distances. Moreover, we focus on the 1-dimensional features, i.e., loops. Further, the hyper-parameter $\sigma^{2}$ in the kernels is taken to be .01. We set $K$ equal to six for illustration purposes. We observe that the TDA-based distance measures AWD and DWD outperform the other distance measures, particularly the existing Wasserstein distance. We also compare the accuracy values reported for the same data set in two recent papers Cheng et al. (2020) and He et al. (2019). The authors proposed a new clustering algorithm and a more accurate distance metric in both papers and compared the performance with different clustering methods. The accuracy values obtained using TDA-based distance measures DWD with affinity propagation clustering and AWD and DWD with K-medoids clustering are higher than the accuracy values reported in Cheng et al. (2020) which is .58 and in He et al. (2019) which is .5886.
6.2 Analyzing the assets selected for index tracking

### v2

@@ 5.3 Strategy 2: Sparse Mean-variance and Global Minimum Variance Portfolios
Panel A in Table 10 shows the out-of-sample performance of the TDA-based sparse mean-variance portfolios. We observe that the portfolio $MV2$ achieves the maximum values for mean, SR, and CEQ, and the minimum value of TR, thus outperforming the other portfolios, including TDA-based portfolios and the benchmark portfolios, namely, $MVA$, the naïve strategy and the benchmark index. For instance, $MV2$ achieves the daily Sharpe ratio of .0733 (.3359 per month / 1.1636 per annum) compared to 0.688 and 0.521 achieved by $MVA$ and $SP500$. Further, the least value of the turnover ratio for $MV2$ indicates less re-balancing across the windows and the formation of stable portfolios. In contrast to the performance of TDA-based index-tracking portfolios, $MV4$ fails to consistently beat other TDA-based portfolios as well as the benchmark portfolios. Nonetheless, these findings confirm our conclusion that the TDA-based selection strategy, in particular, based on kernel $K_{2}$, outperforms the benchmark models, thus providing an excellent asset selection criteria. [^24] We also observe that the $MVA$ portfolio has a higher out-of-sample mean, SR, and CEQ compared to those from the naïve $1/N$ portfolio. We also perform hypothesis tests to assess whether the values from $MVA$ and naïve strategy are statistically significant. We observed that the mean and CEQ from $MVA$ is significantly greater than that from naïve strategy at 99% confidence level while the Sharpe ratios from the two are not significantly different, a similar conclusion was drawn in [25]. Panel B in Table 10 shows the out-of-sample performance of the TDA-based sparse global minimum variance portfolios. A similar conclusion can be drawn that in a $K_{2}$ based portfolio, $GV2$ outperforms other TDA-based portfolios, with the exception that $GV5$ has a lower turnover ratio than $GV2$, but the difference is not very large. Further, $GV2$ beats the benchmark portfolios $GVA$, the naïve, and the index in terms of mean and Sharpe ratio. These findings were further confirmed using the statistical tests for mean, SR, and CEQ; for instance, we observed that $MV2$ has a significantly higher mean, SR, and CEQ than $MVA$ portfolios at a 95% confidence level.
@@ B. Global Minimum Variance Portfolios
COVID-19 Robustness Check Next, we analyze the portfolio performance in the more volatile period from August 2019 to August 2020, which includes the COVID-19 pandemic. Table 12 presents the out-of-sample performance. Panel A clearly shows that the sparse TDA-based portfolio $MV2$ performs really well even in this highly volatile period. More specifically, the TDA-based portfolio $MV_{2}$ outperforms the $MVA$, the naive portfolio, and the benchmark index. This is demonstrated with significant gaps; for instance, the Sharpe ratio of 1.38E-01 achieved by $MV_{2}$ is better and statistically significantly than 1.23E-02 and 1.97E-02 achieved by $MVA$ and the naive portfolio, respectively. Similar to the TDA-based mean-variance portfolios, there is clear pattern that emerges in the performance of TDA-based global minimum variance portfolios. That is, $GV_{2}$ outperform $GVA$ and naïve portfolio, thus confirming our findings that TDA-based portfolios, in particular, based on the kernel $K_{2}$ are not only sparse but also perform superior to traditional models even in turbulent market conditions, safeguarding against significant losses. These results underline the out-of-sample efficacy of TDA-based models in managing risk and delivering returns.
@@
Mean -9.35E-04 4.44E-04 -2.22E-04 2.71E-04 -1.16E-03 -1.98E-04 5.75E-04 5.34E-04 5.96E-04
SR -4.30E-02 1.96E-02 -7.49E-03 1.29E-02 -5.05E-02 -8.29E-03 2.91E-02 1.97E-02 2.43E-02
CEQ -1.17E-03 1.87E-04 -6.60E-04 4.94E-05 -1.42E-03 -4.82E-04 3.80E-04 1.66E-04 2.96E-04
TR 1.40 1.17 1.71 1.47 1.40 1.40
5.4 Sensitivity to Clustering Algorithm
As a robustness check, we replace the APC with two alternative algorithms: $K-$medoids and Hierarchical clustering to demonstrate the effectiveness of TDA based distances, regardless of the employed clustering algorithm.252525 We also tried Density-Based Spatial Clustering of Applications with Noise (DBSCAN); however, given the nature of our data, most points were classified as noise rather than clusters. Therefore, we dropped it from further analysis. To do this, we study the performance of portfolios from Strategy 1 and Strategy 2 over the period September 2020 to August 2022. We use the rolling window approach with a window size of 6 months, shifting by one month, resulting into 18 windows. For $K-$medoids, we obtained the optimal number of clusters by maximizing the average silhouette score across $K=[10,50]$. For consistency, we used the same number of clusters, obtained in $K$-medoids, in Hierarchical clustering. In Hierarchical clustering, we obtained centroids (selected as the point with minimum total distance to other points in the cluster) as the cluster representative for Strategy 2. Table 13 presents the performance of the index tracking portfolios (Strategy 1), while Table 14 present the performance of mean-variance portfolios and global minimum variance portfolios (Strategy 2).
Table 13: Out-of-sample performance of clustering based portfolios $C_{1}-C_{6}$ during September 2020 to August 2022 for different clustering algorithms.
$C_{1}$ $C_{2}$ $C_{3}$ $C_{4}$ $C_{5}$ $C_{6}$
Affinity propagation clustering
TE 5.88E-03 3.59E-03 7.82E-03 2.43E-03 5.38E-03 4.98E-03
EMR -3.76E-04 1.10E-04 -8.33E-05 9.17E-05 3.27E-06 1.15E-05
COR 0.875 0.952 0.748 0.979 0.928 0.937
INFO -6.40E-02 3.07E-02 -1.06E-02 3.77E-02 6.08E-04 2.31E-03
TR 0.344 1.10 1.36 1.46 1.25 0.911
K-Medoids clustering
TE 5.62E-03 3.20E-03 4.83E-03 2.29E-03 4.47E-03 4.24E-03
EMR -4.01E-04 1.89E-04 2.91E-05 1.83E-04 1.93E-04 1.05E-04
COR 0.885 0.962 0.903 0.982 0.947 0.952
INFO -7.14E-02 5.91E-02 6.03E-03 8.01E-02 4.31E-02 2.48E-02
TR 0.679 1.36 2.00 1.68 1.49 1.35
Hierarchical clustering
TE 5.89E-03 4.85E-03 5.98E-03 7.33E-03 5.32E-03 6.01E-03
EMR -2.82E-04 2.30E-05 3.71E-04 -5.21E-04 -6.56E-05 2.46E-04
COR 0.869 0.913 0.850 0.842 0.926 0.909
INFO -4.79E-02 4.75E-03 6.20E-02 -7.10E-02 -1.23E-02 4.10E-02
TR 1.33 2.27 2.45 3.18 2.32 2.29
From Table 13, we observe that the TDA based portfolios emerge as the best-performing portfolios across all clustering methods. In particular, $C_{2}$ and $C_{4}$ are top-performing portfolios in affinity propagation clustering and $K$-medoids clustering, while $C_{3}$ performs well under Hierarchical clustering. For instance, $C_{4}$ attains minimum TE, maximum correlation and maximum information ratio while $C_{2}$ attains the second position in terms of these performance measures in APC and $K$-medoids clustering. The performance of these portfolios is mixed under Hierarchical clustering, with no clear standout. Nonetheless, $C_{3}$ outperforms other algorithms in terms of EMR and INFO. Overall, the findings from Table 13 align with our previous findings that TDA based portfolios outperform other benchmarks. Additionally, while all clustering methods show high correlation (COR) with the benchmark, APC achieves this alignment without excessively high turnover, striking a balance between tracking the benchmark and maintaining low trading costs. The information ratio for APC portfolios is also moderate but stable, indicating a reasonable trade-off between excess return and risk. K-medoids and hierarchical clustering may yield higher INFO values in isolated cases, but their elevated turnover suggests that these risk-adjusted returns may not be sustainable due to the increased transaction costs. From Table 14, we observe that no consistent pattern emerges across clustering algorithms. For instance, MV2 with APC dominates other portfolios, as also observed in our main analysis, while MV6 and MV5 dominates TDA based portfolios with $K$-medoids and hierarchical clustering respectively. In contrast, there is no clear standout performing kernel in GMV portfolios regardless of the clustering algorithm. While APC shows slightly better results for GV4, $K$-medoids works well for GV5 and hierarchical clustering is better for GV1. Similar to Table 13, APC exhibits lower turnover values compared to other clustering algorithms, indicating potentially lower transaction costs. In conclusion, TDA based portfolios outperforms the benchmark across all clustering methods, particularly excelling in affinity propagation clustering.
Table 14: Out-of-sample performance of clustering based portfolios $MV1-MV6$ and $GV1-GV6$ during September 2020 to August 2022 for different clustering algorithms.
Panel A. Mean-Variance Portfolios
MV1 MV2 MV3 MV4 MV5 MV6
Affinity propagation clustering
Mean 2.31E-04 1.84E-03 1.25E-03 8.28E-04 9.94E-05 7.83E-04
SR 1.13E-02 5.84E-02 4.77E-02 3.95E-02 3.93E-03 2.49E-02
CEQ 2.05E-05 1.34E-03 9.07E-04 6.08E-04 -2.20E-04 2.88E-04
TR 1.529 1.974 1.930 1.749 1.573 1.294
K-Medoids clustering
Mean -3.33E-05 4.10E-04 4.78E-04 1.11E-03 8.92E-04 1.34E-03
SR -1.21E-03 1.72E-02 1.74E-02 4.42E-02 3.02E-02 5.38E-02
CEQ -4.10E-04 1.25E-04 1.01E-04 7.95E-04 4.55E-04 1.03E-03
TR 2.359 2.800 3.536 2.337 2.122 2.300
Hierarchical clustering
Mean -1.91E-04 7.79E-05 1.21E-03 1.66E-03 1.77E-03 4.79E-04
SR -5.93E-03 2.63E-03 4.19E-02 5.18E-02 6.59E-02 1.56E-02
CEQ -7.11E-04 -3.59E-04 7.94E-04 1.14E-03 1.41E-03 9.60E-06
TR 2.889 3.458 4.952 3.546 3.261 3.084
Panel B. Global Minimum-Variance Portfolios
GV1 GV2 GV3 GV4 GV5 GV6
Affinity propagation clustering
Mean 1.66E-04 1.43E-04 1.31E-04 4.12E-04 1.60E-04 3.62E-04
SR 1.96E-02 1.64E-02 1.63E-02 4.27E-02 1.01E-02 4.51E-02
CEQ 1.30E-04 1.05E-04 9.85E-05 3.66E-04 1.01E-04 3.30E-04
TR 1.506 1.457 1.822 1.500 1.213 1.226
K-Medoids clustering
Mean 2.08E-04 -1.39E-04 5.39E-04 2.57E-04 6.86E-04 3.38E-04
SR 2.37E-02 -1.49E-02 6.01E-02 2.73E-02 8.23E-02 3.89E-02
CEQ 1.69E-04 -1.83E-04 4.99E-04 2.13E-04 6.52E-04 3.00E-04
TR 2.119 2.280 3.449 2.130 1.934 1.942
Hierarchical clustering
Mean 4.16E-04 -1.41E-04 -1.75E-04 3.86E-04 3.43E-04 3.47E-04
SR 4.98E-02 -1.52E-02 -1.95E-02 4.37E-02 3.89E-02 4.09E-02
CEQ 3.81E-04 -1.84E-04 -2.16E-04 3.47E-04 3.04E-04 3.11E-04
TR 2.681 2.992 4.373 3.346 2.589 2.493
@@ 6 Conclusion
Declaration: There are no financial or non-financial competing interests.
Funding: No funding was received.
Appendix A Affinity Propagation Clustering
The Affinity Propagation (AP) algorithm follows an iterative process that recursively searches for clusters by exchanging real-valued messages between data points, eventually identifying high-quality exemplars that define the resulting clusters. The AP algorithm starts with a similarity measure $s(i,k)$ between sample points $x_{i}$ and $x_{k}$, and begins by treating each data point as a potential candidate for cluster representative, and finds a collection of optimal class representative points based on the preferences, $s(k,k)$ with a larger value of $s(k,k)$ denoting the high likelihood of selecting the sample point $x_{k}$ as class representatives. Initially, AP considers each sample point with equal probability, i.e., giving $s(k,k)$, $\forall k$, the same preference value $P$ [^26] We chose $P$ as the median of the similarity matrix, a common choice in the literature, as it yields a moderate number of clusters. Alternative, selecting the minimum of similarity matrix results in a fewer clusters. The numbers of exemplars are influenced by the preferences as well as the message passing procedure.
To obtain clusters, the AP algorithm primarily exchanges two types of messages: responsibility $r(i,k)$ and availability $a(i,k)$, between sample points, where $r(i,k)$ is sent from data point $x_{i}$ to potential exemplar $x_{k}$ depicting the extent to which $x_{i}$ favors $x_{k}$ as exemplar, and $a(i,k)$ indicates the suitability of selecting $x_{k}$ as a exemplar for $x_{i}$. The algorithm iteratively exchange these messages by updating $r$ and $a$ as follows, eventually determining $k$ as the final cluster center based on the decision matrix $E$,
$r(i,k)\leftarrow s(i,k)-\max_{k^{\prime}\neq k}\{a(i,k^{\prime})+s(i,k^{\prime})\}$
$a(i,k)\leftarrow\begin{cases}\min\left(0,r(k,k)+\sum_{i^{\prime}\neq\{i,k\}}\max(0,r(i^{\prime},k))\right)&\text{if }i\neq k\\ \sum_{i^{\prime}\neq k}\max(0,r(i^{\prime},k))&\text{if }i=k\end{cases}$
where, $a(i,k)$ is set to be zero initially. A damping coefficient, $\lambda\in[0,1]$, is also used to avoid numerical oscillations and ensure stability and speedy convergence such that
$r^{(t+1)}(i,k)\leftarrow(1-\lambda)r^{(t+1)}(i,k)+\lambda r^{(t)}(i,k)$
$a^{(t+1)}(i,k)\leftarrow(1-\lambda)a^{(t+1)}(i,k)+\lambda a^{(t)}(i,k).$
That is, in each iteration, each message takes a weighted combination of its value from previous iteration and the current value, and update all responsibilities given availabilities, availabilities given responsibilities and finally combining them to monitor the exemplar decisions and concluding the AP clustering when these decisions did not change for certain number of iterations or a maximum number of iterations have been executed.
A.1 Case study: Time series clustering
@@
In this section, we demonstrate the superior performance of the newly introduced distance measures by applying clustering on a standard data set from the machine learning literature. More precisely, we show that the distance measures $AWD$ and $DWD$ that quantify the dissimilarity between persistence diagrams ${\mathcal{D}}_{x}$ and ${\mathcal{D}}_{y}$ outperform the conventional Wasserstein distance ($WD$). Furthermore, we show that the proposed distance measure on landscapes $ALD$ and $DLD$ outperform the traditional distance between landscapes ($LD$), defined as $||\lambda_{x}-\lambda_{y}||_{p}$ where $\lambda_{x}$ and $\lambda_{y}$ are landscapes corresponding to the PD ${\mathcal{D}}_{x}$ and ${\mathcal{D}}_{y}$. For comparison purposes, we also compare the traditional distances, denoted by $d_{1}$ and $d_{2}$, defined in terms of Spearman’s rank correlation $\rho_{1}$ and Pearson’s correlation coefficient $\rho_{2}$, as follows $d_{i}(x,y)=\sqrt{2(1-\rho_{i}(x,y))}~{}~{}i\in\{1,2\}.$ Finally, for completeness, we also consider the Euclidean distance $ED$. We evaluate the clustering performance in terms of the accuracy ($ACC$) [40], which ranges from 0 to 1, with a higher value indicating better clustering results.
@@
We collect the data set from the UCI machine learning repository that contains 600 examples of control charts with 60 instances each, synthetically generated by the process in [4]. The data consists of six different classes of control charts: Normal and cyclic, increasing and decreasing trend, upward and downward shift.
@@
Table A.1: Accuracy of clustering results with different distance metrics. Each value $v$ is to be read as $v\times 10^{-2}$
@@
Table A.1 presents the clustering results based on two standard algorithms, the affinity propagation clustering (APC) with $K$ clusters and the $K$-medoids with the distance measures mentioned above. [^27] We use Gaussian exponential kernel as similarity measure defined using these distance measures as an input for APC. We fix embedding dimension $d=2$, and we take $\tau=1$ with $p=1$ for calculating the TDA-based distances. Moreover, we focus on the 1-dimensional features, i.e., loops. Further, the hyper-parameter $\sigma^{2}$ in the kernels is taken to be .01. We set $K$ equal to six for illustration purposes. We observe that the TDA-based distance measures AWD and DWD outperform the other distance measures, particularly the existing Wasserstein distance. We also compare the accuracy values reported for the same data set in two recent papers [19] and [40]. The authors proposed a new clustering algorithm and a more accurate distance metric in both papers and compared the performance with different clustering methods. The accuracy values obtained using TDA-based distance measures DWD with affinity propagation clustering and AWD and DWD with K-medoids clustering are higher than the accuracy values reported in [19] which is .58 and in [40] which is .5886.
Appendix B Sensitivity to window length
We also analyze the sensitivity of the clustering structure to the variations in the size of input window length, across different similarity matrices. We conduct this sensitivity analysis over two consecutive periods: September 2020 to August 2021 and September 2021 to August 2022. The study across two different years allows us to evaluate how changes in the input window length influence the clustering structure and assess the stability of the clustering outcomes across different timeframes. In each period, we construct similarity matrices and hence the clusters for varying input window lengths, starting with one month and expanding upto 12 months. For instance, for the first period, we obtain a similarity matrix using data from the month September 2020. We then extend the input window length incrementally: first to cover two months September 2020 to October 2020, then to three months and finally the period of 12 months, September 2020 to August 2021. We repeat this exercise for the second period, September 2021 to August 2022.
Figure A.1: Similarity between the cluster with benchmark across different window sizes with similarity matrices $K_{1}-K_{6}$ representing the sensitivity of the clustering with respect to the window size (Period - September 2020 to August 2021 ). Note that we have replaced the diagonal entries, which represent self-similarity and are typically 1, with 0 to enhance the contrast, leading to a more visible similarity patterns off-diagonally.
We analyze the evolution of the clustering structure as the window length increases by calculating the adjusted rand index (ARI) between two clustering structures for two different window lengths. We conduct this analysis for all the six kernels, K1 to K6. Resulting heatmaps are presented in Figures A.1 and A.2 for first and second periods respectively. Note that we have replaced the diagonal entries, which represent self-similarity and are typically 1, with 0 to enhance the contrast, leading to a more visible similarity patterns off-diagonally.
Figure A.2: Similarity between the cluster with benchmark across different window sizes with similarity matrices $K_{1}-K_{6}$ representing the sensitivity of the clustering with respect to the window size (Period - September 2021 to August 2022). Note that we have replaced the diagonal entries, which represent self-similarity and are typically 1, with 0 to enhance the contrast, leading to a more visible similarity patterns off-diagonally.
Several observations emerge from these heat maps. First, the kernels $K_{1}$ and $K_{3}$ exhibit occasional patches across both the time-periods, indicating their high sensitivity to the window size and hence relative low stability over different input window lengths. Second, the kernel $K_{4}$ shows moderate similarity across varying window lengths, in particular, when window length is greater than 3 months, suggesting presence of robustness in kernel $K_{4}$. Third, the kernels $K_{2}$, $K_{5}$ and $K_{6}$, show high similarity in adjacent window lengths as well as when the difference in window size is large, signifying their robustness over broader range of the window size. Further, these kernels show two big patches (squares). The kernels $K_{2}$ and $K_{6}$ show similar pattern in the first period (Figure A.1), while kernels $K_{3}$ and $K_{5}$ show similar patterns in the second period (Figure A.2). It is remarkable that kernel $K_{2}$ not only captures what correlation based kernels capture, but also additional features, making it a more comprehensive measure. These findings show that the TDA based kernel $K_{2}$ and $K_{4}$ are more robust to the variation in the window length across different time-periods in contrast to correlation based distance measures that show period-dependent behavior.
Appendix C Analyzing the assets selected for index tracking